Arthur bought three hamburgers and four hot dogs for $10. The next day, he bought two hamburgers and three hot dogs for $7. How much does a hot dog cost?
Answer by mananth(16949)

hamburgers x cost
hotdogs ------------ y cost
3x+4y=10.............1
2x+3y=7 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 4
-9x-12y =-30
8x+12y=28
Add the two equations
-1x=-2
/-1
x=$2 hamburgers
plug value of x in (1)
3x+4y= 10
6+4y=10
4y=10-6
4y=4
y=$1 hotdogs